OpenStudy (anonymous):

how would you factor 3y^2-7y+2

jhonyy9 (jhonyy9):

y_1,-2=(7+/- sqrt(49-24))/6 =(7+/- 5)/6 = 2/6 = 1/3 and 12/6 = 2 so than factored will be (3y-1)(y-2)